# One Wave to Explain Them All: A Unifying Perspective on Post-hoc Explainability

### v1

@@ 3 Methods
which can be rewritten as a convolution (Mallat, 2008). In the discrete dyadic case, the scale factor $\bm{\lambda}$ takes values in a set $\bm{\Lambda}$, chosen as $\bm{\Lambda}=\{2^{j}:1\leq j\leq N\;,N\in\mathbb{N},N>0\}$. Mallat (1989) showed that one can compute the dyadic wavelet transform of a signal $\bm{g}$ by applying a high-pass filter $H$ to the signal $\bm{g}$ and subsampling by a factor of two to retrieve the detail coefficients, and applying a low-pass filter $G$ and subsampling by a factor of two to retrieve the approximation coefficients. Iterating on the approximation coefficients generates a multilevel transform, where the $j^{th}$ level extracts information at resolutions between $2^{j}$ and $2^{j-1}$ octaves in the frequency spectrum. When the input signal $\bm{x}$ has dimensionality greater than one, its detail coefficients can be decomposed into different orientations. The common orientations for 2D signals (i.e., images) are vertical, horizontal, and diagonal.
Wavelets and multiscale decompositions.
Multiscale analysis consists in decomposing an input signal into different levels of detail. The resulting decomposition is particularly interesting as it generates interesting features for signal understanding: edges in images at different orientations and scales correspond to different textures. In sounds, the multiscale decomposition isolates slowly changing patterns from transient ones. Overall, the wavelet decomposition enables the decomposition of an input signal into interpretable components. As we further discuss in section 4.2, the properties of multiscale decompositions translate into several insightful properties for XAI.
3.1 Gradient-based feature attribution in the wavelet domain
@@
Let $\bm{f}$ be a classifier and $\bm{x}$ an input (e.g., an image, an audio, or a 3D shape). The classifier $\bm{f}$ maps the input to a class $c$ as $\bm{y}_{c}=\arg\max_{c\in\mathcal{C}}\bm{f}(\bm{x})\equiv\bm{f}_{c}(\bm{x})$ with a slight abuse of notation. We recall that the original saliency map of the classifier $\bm{f}$ for class $c$ is then given by $\bm{\gamma}_{\text{Sa}}(\bm{x})=|\nabla_{\bm{x}}\bm{f}_{c}(\bm{x})|$ where $c$ denotes the class of interest. The saliency map is defined provided that the $\bm{f}_{c}$’s are piecewise differentiable (Simonyan et al., 2014). The saliency map highlights the most influential (in terms of the absolute value of the gradient) components in the input $\bm{x}$ for determining the model’s $\bm{f}$ decision. The higher the value, the greater the importance of the corresponding region.
However, varying pixel values provide no information to what is changing on the image. Therefore, we argue that the pixel domain is not well suited for explaining what the model is seeing on the image. On the other hand, the wavelet decomposition of an image – and more broadly of any differentiable modality – provides information on the structural components of the modality. Therefore, computing the gradient of $\bm{f}$ with respect to the wavelet transform of $\bm{x}$ will enable us to understand the model’s reliance on features such as textures, edges, or shapes in the case of images, transients, or harmonics in sounds or corners or small details in 3D shapes.
Denoting $\bm{z}=\mathcal{W}(\bm{x})$ the wavelet transform of $\bm{x}$, since $\mathcal{W}$ is invertible, we can define the saliency map of in the wavelet domain as
@@
using the fact that $\bm{x}=\mathcal{W}^{-1}(\bm{z})$ and where ${\frac{\partial\bm{f}_{c}(x)}{\partial\bm{x}}}$ denotes the gradient of the classifier output with respect to the input image and ${\frac{\mathcal{W}^{-1}(\bm{z})}{\partial\bm{z}}}$ is the Jacobian matrix of the inverse wavelet transform. In practice, to retrieve Equation 1, we require the gradients on $\mathcal{W}(\bm{x})$ and directly evaluate ${\partial\bm{f}_{c}(\mathcal{W}^{-1}(\bm{z}))/\partial\bm{z}}$. A remarkable property of this framework is that it accommodates any input dimension, and thus it is modality-agnostic. Therefore, we can apply it – and leverage its properties – to numerical signals such as audio (1D signals), images (2D signals), or 3D shapes (3D signals). In this paper, we demonstrate the superiority of this method in the 1D and 2D settings compared to other domain-specific methods and illustrate examples for 3D classification.
@@
Smilkov et al. (2017) highlighted the fact that the saliency maps computed following Equation 1 can fluctuate sharply at small scales as $\bm{f}_{c}$ is not continuously differentiable. To yield smoother explanations, Smilkov et al. (2017) perturb the input image with Gaussian noise. Analogously, we propose to calculate
$\bm{\gamma}_{\text{Sg}}(\bm{z})=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\tilde{\bm{z}}}\bm{f}(\mathcal{W}^{-1}(\tilde{\bm{z}}))~{}~{}~{}\text{with}~{}~{}~{}\tilde{\bm{z}}=\mathcal{W}(\bm{x}+\bm{\delta})~{}~{}\text{and}~{}~{}\bm{\delta}\sim\mathcal{N}(0,I\sigma^{2}).$ (2)
The number of samples $n$ needed to compute the approximation of the smoothed gradient and the standard deviation $\sigma^{2}$ are hyperparameters for their method. To transpose this method to the wavelet domain, we add noise to the input before computing its wavelet transform. We refer to this method as WAMSG throughout the rest of the paper. In appendix A.1, we illustrate the enhancement of the quality of the explanation after applying the smoothing to the gradients as described in equation 2.
@@
Another approach to derive smooth explanations from the model’s gradients consists in averaging the gradient values along the path from a baseline state to the current value. The baseline state is often set to zero, representing the complete absence of features. This technique, introduced by Sundararajan et al. (2017), satisfies two axioms, sensitivity and implementation invariance. Sensitivity states that “for every input and baseline that differ in one feature but have different predictions, then the differing feature should be given a non-zero attribution” and Implementation Invariance that “the attributions are always identical for two functionally equivalent networks”. Following Sundararajan et al. (2017), we adapt the Integrated Gradient method from the image domain to the wavelet domain. Denoting $\bm{z}=\mathcal{W}(\bm{x})$, we evaluate
@@
$\bm{\gamma}_{\text{Ig}}=(\bm{z}-\bm{z}_{0})\cdot\int_{0}^{1}\frac{\partial\bm{f}_{c}\left(\mathcal{W}^{-1}\left(\bm{z}_{0}+\alpha(\bm{z}-\bm{z}_{0})\right)\right)}{\partial\bm{z}}\mathrm{d}\alpha,$ (3)
where $\bm{z}_{0}$ denotes the baseline state of the wavelet decomposition of $\bm{x}$. We refer to this implementation of WAM as WAMIG. In appendix A.1, we illustrate the enhancement of the quality of the explanation after applying the smoothing to the gradients as described in Equation 2 and after integrating the gradients, as described in Equation 3. We also discuss the visualization properties that emerge when using either method.
3.2 Modality-specific declinations
Images.
For images, the computation of WAM following Equation 2 can be visualized on the dyadic wavelet transform of the image (see plot (b) of Figure 2. From the wavelet transform of the image, we can derive several interesting properties regarding what the model sees on the input image. In particular, we can decompose the important coefficients at each scale (d), and illustrate as a reconstructed image (c) what is important for the model’s prediction. In section 4.2, we discuss how WAM enables us to revisit the meaningful perturbation framework of Fong & Vedaldi (2017) and introduce minimal images, and in appendix C, we show the connections between our method and the frequency-centric perspectives on model robustness of Zhang et al. (2022) or Yin et al. (2019).
Audio.
Following Equation 2 and Equation 3, we obtain the sensitivity of the model’s prediction with respect to the wavelet coefficients of the waveform. As in practice, audio classifiers take mel-spectrograms as inputs; we also retrieve the gradients with respect to the mel-spectrogram of the waveform. Therefore, our approach bridges the gap between methods that focus on explanations of the waveforms and those that return explanations on the mel-spectrogram.
3D Shapes.
We apply WAM on voxels, as the wavelet transform is defined on structured data formats. Voxels represent 3D space as a grid of uniformly spaced cubes, where each voxel stores information about the object at that position. We consider the 3D wavelet transform, which generalizes the wavelet transform to upper dimensions. Implementations of WAMSG and WAMIG follow Equation 2 and Equation 3 analogously to images and audios.
Figure 2: WAM for images. Our method decomposes the important components at different scales (i.e., different levels of details) and enables us to see what is seen on the image by the model.
3.3 Evaluation
We evaluate WAM in two distinct settings: images and audio. Evaluation is carried out on usual benchmarks for both modalities. We do not consider the 3D setting for quantitative evaluation due to the lack of comparable baselines.
Common evaluation metrics.
@@ 3.3 Evaluation
We quantitatively assess the accuracy of our method by leveraging the Faithfulness (Muzellec et al., 2023), defined as the difference between the Insertion and the Deletion scores, introduced by Petsiuk et al. (2018). Insertion and Deletion have been widely used in XAI to evaluate the quality of feature attribution methods (Fong & Vedaldi, 2017). The Deletion measures the evolution of the prediction probability when one incrementally removes features by replacing them with a baseline value according to their attribution score. Insertion consists in gradually inserting features into a baseline input. Samek et al. (2016) and Li et al. (2022) have shown that Faithfulness is effective in evaluating attribution methods. Given a model $\bm{f}$ and an explanation functional $\bm{\gamma}$, the Faithfulness $F$ is given by
@@
We provide a detailed derivation of the Insertion and the Deletion scores in appendix B.1. Insertion and Deletion were initially defined in the context of images, but we propose a definition that expands them to audio.
Modality-specific metrics.
In addition to the Faithfulness, we compare WAM using the $\mu$-Fidelity (Bhatt et al., 2021) for images and the Faithfulness on Spectra (FF, Parekh et al., 2022) and Input Fidelity (Fid-In, Paissan et al., 2023) for audio. We refer the reader to the appendix B.2 for a thorough definition of these metrics and a discussion of the results.
Evaluation setting for images.
For images, we evaluate our method on a subset of the validation set of ImageNet (Russakovsky et al., 2015). Our subset contains 1,000 images randomly sampled from the 50,000 images of the validation set of ImageNet. We consider four model architectures representative of the popular topologies currently used. We consider the following models: the ResNet (He et al., 2016), the ConvNext (Liu et al., 2022), the EfficientNet (Tan & Le, 2019) and the Data efficient transformer (DeiT, Touvron et al., 2021). We refer the reader to the appendix A.2 for more details on the model’s parametrizations that we used. This evaluation framework is based on the frameworks of Fel et al. (2021).
We compare our method with alternative gradient-based methods, namely Saliency (Simonyan et al., 2014), Integrated Gradients (Sundararajan et al., 2017), GradCAM and GradCAM++ (Selvaraju et al., 2017), and SmoothGrad Smilkov et al. (2017). We focus only on gradient-based methods, as they are more faithful and faster to generate than alternative approaches (Crabbé & van der Schaar, 2023; Wang & Wang, 2021; Xue et al., 2023).
Evaluation setting for audio.
We evaluate our method on the dataset for Environmental Sound Classification (ESC-50, Piczak, 2015). We pick the 400 samples of the first fold of ESC-50, as our backbone model has been trained on the remainder of the dataset, and evaluate the CNN classification model of Kumar et al. (2018) as our black-box model to explain. We consider a single model as alternative models (Huang & Leanos, 2018; Wilkinghoff, 2021; Lopez-Meyer et al., 2021) are only variations around the same topology. We consider two variants of the ESC-50 dataset: the original (unaltered samples) and noisy samples, for which we add 0 dB white noise to the input samples whose prediction we seek to explain. We include three baseline methods, which return explanations on the mel-spectrogram of the input samples: the GradCAM (Selvaraju et al., 2017), Integrated Gradients (Sundararajan et al., 2017), SmoothGrad (Smilkov et al., 2017) and Saliency (Simonyan et al., 2014).
4 Results
4.1 Quantitative evaluation results
Images.
As displayed on Table 1, we can see that WAM for 2D signals outperforms competing baselines according to the Faithfulness metric. In appendix B.2, we present additional results using the Insertion, Deletion, and the $\mu$-Fidelity. The good results are mostly driven by the fact that WAM performs well in terms of Insertion. WAM also passes the randomization test (Adebayo et al., 2018). We refer the reader to appendix B.3 for more details on this test.
Table 1: Faithfulness (Muzellec et al., 2023) score obtained on 1,000 images from the validation set of ImageNet and for different model architectures. Higher is better. Bolded results are the best and underlined values are the second best.
Method ResNet ConvNext EfficientNet DeiT Mean
Saliency 0.025 0.032 0.008 0.038 0.025
Integrated Gradients 0.000 0.001 0.000 0.003 0.001
GradCAM 0.134 0.072 0.061 0.162 0.107
GradCAM++ 0.184 0.055 0.050 0.044 0.083
SmoothGrad 0.023 0.000 0.010 0.004 0.009
WAMSG (ours) 0.438 0.334 0.350 0.423 0.386
WAMIG (ours) 0.344 0.359 0.370 0.420 0.373
Audio.
Table 2 presents the evaluation results for audio. For WAM, we generate the explanations from the wavelet coefficients. We can see that for audios, WAM also achieves state-of-the-art results and outperforms the competing metrics in terms of Faithfulness of spectra, Input fidelity and Insertion. The results of the other metrics are in line with those of competing approaches. In appendix B.2, we provide additional results where explanations are computed from the mel-spectrogram of the waveform. In this case, we report that WAM’s performance is more in line with competing approaches, thus showing the added value brought by explaining the model’s decision through the wavelet domain.
Table 2: Evaluation scores of WAM and comparison with baselines on 400 audio samples from ESC-50 (fold 1). The column ”ESC” indicates that the samples are unaltered. The column ”+WN” indicates that the samples have 0 dB Gaussian white noise. We report the results with explanations generated from the wavelet coefficients of the waveform. Bolded results are the best and underlined values are the second best.
Method Faithfulness ($\uparrow$) Insertion ($\uparrow$) Deletion($\downarrow$) FF ($\uparrow$) Fid-In ($\uparrow$)
ESC50 +WN ESC50 +WN ESC50 +WN ESC50 +WN ESC50 +WN
IntegratedGradients 0.264 0.310 0.267 0.312 0.047 0.045 0.207 0.207 0.220 0.225
GradCAM 0.072 0.073 0.274 0.274 0.201 0.201 0.137 0.135 0.517 0.542
Saliency 0.066 0.065 0.220 0.221 0.154 0.156 0.166 0.168 0.253 0.245
WAMSG (ours) 0.197 0.205 0.449 0.452 0.252 0.246 0.132 0.130 0.718 0.690
WAMIG (ours) 0.176 0.182 0.436 0.442 0.260 0.261 0.118 0.124 0.652 0.657
4.2 XAI Properties of the Wavelet Attribution Method
Revisiting Meaningful Perturbation.
In the seminal works by Fong & Vedaldi (2017) and Fong et al. (2019), a method is introduced to explain the most important parts of an image by optimizing a mask $\bm{m}$ that partially occludes certain regions. The objective is twofold: (i) preserve the classification score, while (ii) remove as much of the image content as possible to isolate the most relevant features. However, optimizing in the pixel domain presents challenges in producing smooth masks (Fong et al., 2019), necessitating various regularization techniques, smoothing operations, and data augmentations to mitigate these issues. We revisit this framework by proposing to recast the problem in the wavelet domain as a more suitable space for optimization. The wavelet domain inherently captures spatial and spectral information, providing a natural structure for producing meaningful and interpretable solutions. Specifically, we solve the following optimization problem:
$\bm{m}^{\star}=\operatorname*{arg\,min}_{\bm{m}\in[0,1]^{|\mathcal{X}|}}\bm{f}_{c}\left(\mathcal{W}^{-1}(\bm{z}\odot\bm{m})\right)+\alpha\|\bm{m}\|_{1},$
$\bm{m}_{i+1}=\bm{m}_{i}-\eta\nabla_{\bm{m}_{i}}\left(\bm{f}_{c}\left(\mathcal{W}^{-1}(\bm{z}\odot\bm{m}_{i})\right)+\alpha\|\bm{m}_{i}\|_{1}\right),$
From an interpretability perspective, our method offers significant and novel insights. Traditional meaningful perturbation methods (Fong & Vedaldi, 2017) focus on spatial localization, identifying clusters of pixels that answer the question of where the important features are located. However, this spatial emphasis alone provides a limited understanding of the underlying data structure. In contrast, by operating in the wavelet domain, our method captures both the what – the relevant scales – and the where – their spatial locations. This dual information enriches the explanation by revealing the location and the nature of the features influencing the model’s decision.
Figure 3: A) Sparsity-optimized minimal images. We revisit meaningful perturbation by optimizing the sparsity of the wavelet transform using masking, instead of optimizing the mask in pixel space. The displayed examples show that the resulting minimal images reveal the model’s reliance on textures. B) Sparsity Pareto front. As $\alpha$ increases, the sparsity of the wavelet coefficients increases (x-axis), but beyond a certain point, too much information is lost and the logit score drops to zero. However, we observe that many components can be removed before adversely affecting the model. Results are averaged across 1,000 images optimized for 500 steps, and for $\alpha$ ranging in $[0,100]$ for each image.
By leveraging the wavelet domain, our approach addresses the challenges associated with pixel-space optimization and provides a more comprehensive understanding of the model’s behavior. This method generalizes across different data modalities and can be a valuable tool for interpreting complex neural networks in various applications.
Audio: post-hoc identification of relevant parts of the input audio.
Figure 4 qualitatively illustrates an application of WAM for audio signals. Herein, we perform a noise experiment to add 0 dB white noise to a target audio to form the input audio. The models prediction does not alter after introducing the noise and thus the model is expected to still rely on parts of input audio coming from target audio for its decision. The interpretation audio in Figure 14 generated using top wavelet coefficients provides insights into the decision process and supports this hypothesis. In particular, it almost entirely filters out corruption audio and without requiring any training it also clearly emphasizes key parts of target audio. Similarly, we discuss in appendix C how WAM also retrieves the key parts of an audio signal that has been corrupted with another source (overlap experiment).

### v2

@@ 3.1 Multiscale decompositions
Mallat (1989) showed that the dyadic wavelet transform of a signal $\bm{g}$ can be computed by applying a high-pass filter $h$, followed by downsampling by a factor of two, to obtain the detail coefficients, and similarly applying a low-pass filter $g$, followed by downsampling, to obtain the approximation coefficients. Iteratively applying this to the approximation coefficients yields a multilevel transform, where the $j^{\text{th}}$ level captures information at scales corresponding to frequency bands between $2^{j}$ and $2^{j-1}$ octaves. For input signals $\bm{g}$ of dimension greater than one, the detail coefficients can be decomposed into directional components. In the 2D case (e.g., images), these typically correspond to vertical, horizontal, and diagonal orientations.
3.2 Feature attribution in the wavelet domain
@@
Let $\bm{f}$ be a classifier and $\bm{x}$ an input (e.g., an image, an audio, or a volume). The classifier $\bm{f}$ maps the input to a class $c$ as ${\bm{y}_{c}=\arg\max_{c\in\mathcal{C}}\bm{f}(\bm{x})\equiv\bm{f}_{c}(\bm{x})}$ with a slight abuse of notation. We recall that the original saliency map of the classifier $\bm{f}$ for class $c$ is then given by $\bm{\gamma}_{\text{Sa}}(\bm{x})=|\nabla_{\bm{x}}\bm{f}_{c}(\bm{x})|$ where $c$ denotes the class of interest. The saliency map is defined under the condition that the $\bm{f}_{c}$’s are piecewise differentiable (Simonyan et al., 2014). It highlights the most influential (in terms of the absolute value of the gradient) components in the input $\bm{x}$ for determining the model’s $\bm{f}$ decision. The higher the value, the greater the importance of the corresponding region.
However, moving from one pixel to the next is only a shift in the spatial domain and does not capture relationships between scales or frequencies. Therefore, we argue that the pixel domain is not well suited for explaining what the model is seeing on the image. On the other hand, the wavelet decomposition of an image – and more broadly of any differentiable modality – provides information on the structural components of the modality. Therefore, computing the gradient of $\bm{f}$ with respect to the wavelet transform of $\bm{x}$ will enable us to understand the model’s reliance on features such as textures, edges, or shapes in the case of images and volumes and transients, or harmonics in sounds.
Denoting $\bm{z}=\mathcal{W}(\bm{x})$ the discrete wavelet transform of $\bm{x}$, since $\mathcal{W}$ is invertible, we can define the saliency map in the wavelet domain as
@@
using the fact that $\bm{x}=\mathcal{W}^{-1}(\bm{z})$ and where ${\frac{\partial\bm{f}_{c}(x)}{\partial\bm{x}}}$ denotes the gradient of the classifier output with respect to the input image and ${\frac{\mathcal{W}^{-1}(\bm{z})}{\partial\bm{z}}}$ is the Jacobian matrix of the inverse wavelet transform. In practice, to retrieve Equation 6, we require the gradients on $\mathcal{W}(\bm{x})$ and directly evaluate ${\frac{\partial\bm{f}_{c}(\mathcal{W}^{-1}(\bm{z}))}{\partial\bm{z}}}$.
A remarkable property of this framework is that it accommodates any input dimension, and thus is modality-agnostic. Therefore, we can apply it – and leverage its properties – to numerical signals such as audio (1D signals), images (2D signals), or volumes (3D signals). In this paper, we demonstrate the superiority of this method in the 1D, 2D and 3D settings compared to other domain-specific methods.
@@
Saliency maps computed following Equation 6 can fluctuate sharply at small scales as $\bm{f}_{c}$ is not continuously differentiable. To yield smoother explanations, smoothing consists in averaging the explanation over a set of noisy samples (Smilkov et al., 2017). Analogously, we propose to calculate
$\bm{\gamma}_{\text{SG}}(\bm{z})=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\tilde{\bm{z}}}\bm{f}(\mathcal{W}^{-1}(\tilde{\bm{z}})),$ (7)
where $\tilde{\bm{z}}=\mathcal{W}(\bm{x}+\bm{\delta})$ and $\bm{\delta}\sim\mathcal{N}(0,I\sigma^{2})$. The number of samples, $n$, needed to compute the approximation of the smoothed gradient and the standard deviation $\sigma^{2}$ are hyperparameters. To transpose this method to the wavelet domain, we add noise to the input before computing its wavelet transform. We refer to this method as WAMSG throughout the rest of the paper. In Section A.1, we illustrate the enhancement of the quality of the explanation after applying the smoothing to the gradients as described in Equation 7.
@@
Another approach to derive smooth explanations from the model’s gradients consists in averaging the gradient values along the path from a baseline state to the current value. The baseline state is often set to zero, representing the complete absence of features. This technique, Integrated Gradients (Sundararajan et al., 2017), satisfies two axioms, sensitivity and implementation invariance. Sensitivity states that “for every input and baseline that differ in one feature but have different predictions, then the differing feature should be given a non-zero attribution” and implementation invariance that “the attributions are always identical for two functionally equivalent networks”. We adapt the Integrated Gradient method from the image domain to the wavelet domain. Denoting $\bm{z}=\mathcal{W}(\bm{x})$, we evaluate
@@
$\bm{\gamma}_{\text{IG}}=(\bm{z}-\bm{z}_{0})\cdot\\ \int_{0}^{1}\frac{\partial\bm{f}_{c}\left(\mathcal{W}^{-1}\left(\bm{z}_{0}+\alpha(\bm{z}-\bm{z}_{0})\right)\right)}{\partial\bm{z}}\mathrm{d}\alpha,$ (8)
where $\bm{z}_{0}$ denotes the baseline state of the wavelet decomposition of $\bm{x}$. We refer to this implementation of WAM as WAMIG.
Section A.1 illustrates the enhancement of the quality of the explanation after applying the smoothing to the gradients as described in Equation 7 and after integrating the gradients, as described in Equation 8.
Theoretical properties.
WAM extends feature attribution into the wavelet domain. The main focus of the paper is the proposal and the practical usage of WAM, however, it appears that WAMIG inherits from the theoretical properties of the vanilla integrated gradients method, and especially sensitivity. We refer the reader to Appendix B for a discussion on the theoretical aspects of WAM.
4 Evaluation and applications
4.1 Quantitative evaluation
Evaluation metrics.
@@ 4.1 Quantitative evaluation
We evaluate WAM against gradient-based methods, as they are more reliable and efficient than perturbation-based methods (Crabbé & van der Schaar, 2023). We evaluate the Faithfulness, which we define as the difference between the Insertion and Deletion scores, thereby following the definition proposed by Muzellec et al. (2024). We refer the reader to Chan et al. (2022) for a discussion of the alternative definitions of Faithfulness. Faithfulness is effective in evaluating attribution methods (Samek et al., 2016; Li et al., 2022). Insertion and deletion (Petsiuk et al., 2018) have been widely used in XAI to evaluate the quality of feature attribution methods (Fong & Vedaldi, 2017). The deletion measures the evolution of the prediction probability when one incrementally removes features by replacing them with a baseline value according to their attribution score. Insertion consists in gradually inserting features into a baseline input.
Given a model $\bm{f}$ and an explanation functional $\bm{\gamma}$, the Faithfulness $F$ is given by
@@
We provide a detailed derivation of the Insertion and the Deletion scores in Section C.2. They were initially defined in the context of images, but we propose to expand them to audio and volumes.
Evaluation setting.
To the best of our knowledge, no prior work has evaluated a feature attribution method across multiple input modalities. To address this gap, we designed a comprehensive evaluation framework spanning diverse datasets: ESC-50 (Piczak, 2015) for audio, ImageNet (Russakovsky et al., 2015) for images, and MedMNIST3D (Yang et al., 2023) for volumes. For each modality, we picked an arbitrary model and applied the same feature attribution methods to ensure consistency. We considered four feature attribution methods: SmoothGrad (Smilkov et al., 2017), Saliency (Simonyan et al., 2014), Integrated Gradients (Sundararajan et al., 2017) and GradCAM (Selvaraju et al., 2017). We use the Python library Captum (Kokhlikyan et al., 2020) for consistently implementing existing methods on our datasets for audio and images. For volumes, we integrated WAM as an additional method within the evaluation framework LATEC (Klein et al., 2024).
Results.
Table 1 presents the performance of WAM. Our method consistently outperforms traditional attribution methods, especially for images and volumes, while achieving competitive results for audio. We can see that WAMIG outperforms WAMSG. As further discussed in Section A.1, this improved performance arises because path integration captures inter-scale dependencies, revealing the relative importance of each scale in the model’s prediction. Overall, these results highlight the relevance of using wavelets, rather than pixels, as the basis for feature attribution.
In Section C.3, we evaluate WAM using additional datasets, metrics, and alternative model topologies, and compare its accuracy with more methods across all three input modalities. Our results further emphasize that WAM consistently matches or surpasses the performance of existing methods.
Table 1: Faithfulness (Faith), Insertion (Ins) and Deletion (Del) scores across modalities. Evaluations are conducted on 400 samples from ESC-50 (fold 1), 1,000 images from ImageNet’s validation set and the full AdrenalMNIST3D test set (298 samples). WAM outperforms all other methods on image and volume modalities for both Insertion and Deletion metrics, meaning that it is good at identifying which features influence the most the model’s decision. For audio, it achieves state-of-the-art performance in Insertion and performs as well as other methods for Deletion and Faithfulness. Best results are bolded and second best underlined.
Audio Images Volumes
Model ResNet EfficientNet 3D Former
Dataset ESC-50 ImageNet AdrenalMNIST3D
Ins ($\uparrow$) Del ($\downarrow$) Faith ($\uparrow$) Ins($\uparrow$) Del ($\downarrow$) Faith ($\uparrow$) Ins ($\uparrow$) Del ($\downarrow$) Faith ($\uparrow$)
Integrated Gradients 0.267 0.047 0.264 0.113 0.113 0.000 0.666 0.743 -0.077
SmoothGrad 0.251 0.067 0.184 0.129 0.119 0.010 0.680 0.731 -0.051
GradCAM 0.274 0.201 0.072 0.364 0.303 0.061 0.689 0.744 -0.055
Saliency 0.220 0.154 0.066 0.148 0.140 0.008 0.751 0.742 0.009
WAMIG (ours) 0.436 0.260 0.176 0.447 0.049 0.370 0.719 0.621 0.098
WAMSG (ours) 0.449 0.252 0.197 0.419 0.097 0.350 0.718 0.648 0.070
4.2 Exploring how WAM renews feature attribution
Beyond its strong quantitative performance, WAM also offers new qualitative insights by disentangling features at different scales. This section highlights three representative use cases. We provide more examples in Appendix D.
Images: bridging attribution and robustness.
In computer vision, several works document a correlation between the reliance on low frequency components of the image and the robustness of the model to adversarial or natural corruptions (Zhang et al., 2022; Chen et al., 2022; Wang et al., 2020). This property is grounded in the frequency or spectral bias of models (Rahaman et al., 2019) and has been highlighted by filtering out frequency bands from the Fourier transform of the input images. As the Fourier magnitude discards spatial information, frequencies are removed in all the image.
WAM quantifies the importance of each scale in the final prediction. Scales in the wavelet domain correspond to dyadic frequency ranges in the Fourier domain. Therefore, by summing the importance of wavelet coefficients at a given scale, we can deduce the importance of the corresponding frequency range in the model’s prediction. This estimation can be straightforwardly derived from the explanation and does not require multiple perturbations as done in previous works (Zhang et al., 2022; Chen et al., 2022). As a result, WAM can be used to evaluate the robustness of a model.
To show this property, we compare a set of ResNet models. The baseline is trained using the standard empirical risk minimization (Vanilla ResNet). The others are trained using three adversarial training methods: ADV (Madry et al., 2018), ADV-Fast (Wong et al., 2020), ADV-Free (Shafahi et al., 2019).
Figure 3 averages the importance of each scale for 1,000 explanations computed from ImageNet. For each image, we normalize the importance to highlight the relative reliance on different scales. As expected, the vanilla ResNet relies more on fine scales (high frequencies) than the adversarially trained models. WAM consistently retrieves existing results from the robustness literature while avoiding the need for complex experimental settings, thus paving the way for easier assessment of the robustness of a model.
Figure 3: Model robustness assessment with WAMIG. Each bar indicates the importance of each scale in the model’s prediction. Importances are averaged over 1,000 explanations computed from ImageNet and normalized. The adversarially robust models rely more on coarse-scale (low frequency) features than the vanilla ResNet, showing that WAM recovers results from the literature.
Volumes: revealing multi-scale structures.
We retrieve on voxels the same decomposition as for images or audio. Figure 4 highlights the significance of the edges at larger scales, whereas, at smaller scales, the importance becomes increasingly concentrated at the center of the digit. This multi-scale decomposition is particularly valuable for volumetric data, as it allows us to disentangle fine-grained details from broader structural patterns, which is crucial in many applications. For instance, in medical imaging, larger-scale features may correspond to organ boundaries or lesion contours, while finer scales capture subtle textural variations indicative of disease. Similarly, in 3D object recognition, coarse scales help identify overall shapes, whereas finer scales refine the details essential for distinguishing between similar objects. To the best of our knowledge, WAM is the first method to demonstrate such a decomposition on shapes, offering a new perspective on how models process hierarchical spatial information in 3D data.
Figure 4: Multi-scale decomposition of feature importance on a volume using WAM. Coarse scales (yellow) highlight the edges of the number, capturing its global structure. Fine scales (blue) focus on the digit’s center, capturing high-frequency details and localized variations.
Audio: identifying key sound components.
Figure 5 qualitatively illustrates an application of WAM for audio signals. We perform a noise experiment by adding 0 dB white noise to a target audio to form the input audio. The model’s prediction remains unchanged after adding noise, indicating it continues to focus on the relevant parts of the input audio. Identifying key parts of the audio input is crucial for improving model interpretability and ensuring it focuses on meaningful signal components. Traditionally, this task has been addressed using trained models such as Non-Negative Matrix Factorization (NMF, see for instance Parekh et al., 2022), which requires training to extract relevant features and explain decisions.
In contrast, WAM performs this identification in a post-hoc manner, simplifying the task by directly highlighting the important audio components without the need for prior training. The interpretation audio in Figure 5, generated using top wavelet coefficients, provides further insights into the decision process and is consistent with the fact that the model has not been affected by the addition of noise. Therefore, WAM effectively filters out the corrupted audio while highlighting the key parts of the audio. Similarly, we discuss in Section D.3 how WAM also retrieves the key parts of an audio signal that has been corrupted by another source (overlap experiment).